[Ncep.nhc.nco_contacts] TSB Morning Rounds - Fri Jan 13, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Fri Jan 13 13:30:21 UTC 2017


Here is today's summary of NHC computer operations:
--- Mark

----------

TAFB:

HSU:
The Himawari data are not coming in to N-AWIPS. This has been out for a few  
weeks. Monica is investigating.
Other items:
The major AWIPS update this week was successful. TSB staff, with help from  
Jeff Lewitsky and NCF, are still cleaning up some lingering problems that  
resulted from the new builds.

Some AC issues were reported by TAFB and TSB yesterday. TSB will refer  
those to Jann and William.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170113/8503d147/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list